# Complexity of deep computations via topology of function spaces

### v2

@@ A Compositional Computation Structure (CCS) is a triple $(L,\mathcal{P},\Gamma)$, where
For a CCS $(L,\mathcal{P},\Gamma)$, we regard the elements of $\Gamma$ as “standard” finitary computations, and the elements of $\overline{\Gamma}$, i.e., deep computations, as possibly infinitary limits of standard computations. The main goal of this paper is to study the computability, definability and computational complexity of deep computations. Since ultracomputations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
@@ 2.2. Computability and definability of deep computations and the Extendibility Axiom
    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{L}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
@@
    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model the theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]). The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
@@
To summarize, a function $f:L\to\mathcal{L}$ is computable if and only if it is definable if and only if it is polynomially approximable if and only it can be extended to a continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$. This motivates the following definition.
@@
By the preceding remarks, the Extendibility Axiom says that the elements of the semigroup $\Gamma$ are finitary computations. For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$. For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.
@@ 2.3. Newton’s method as a CCS
Let $p(z)$ be a non-constant polynomial with complex coefficients. Newton’s method is an iterative method that is used to approximate a root of $p(z)$. Define the Newton map as:
@@
The method consists of taking an initial guess $z_{0}\in\mathbb{C}$ and iterating the rational map $N_{p}$ to obtain a sequence given by
@@
For each root $r\in\mathbb{C}$ of $p(z)$, there exists $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
Denote by $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ (nth iteration of $N_{p}$). Given a root $r$ of $p(z)$, the set $B_{r}=\{z\in\mathbb{C}:\lim_{n\rightarrow\infty}N_{p}^{n}(z)=r\}$ is an open set called the basin of $r$. However, Newton’s method can fail to converge to any root for some choices of $z_{0}$. For example, consider the polynomial $p(z)=z^{3}-2z+2$. The Newton map is given by
@@
Notice that taking $z_{0}=0$ as an initial guess will yield the sequence $0,1,0,1,0,1,\dots$ that oscillates between 0 and $1$ but none of them are roots of $p(z)$. Another more chaotic way Newton’s method can fail to converge is when the sequence of iterations has no convergent subsequence. The set of such points is known as the Julia set associated to $N_{p}$ and it is typically a fractal. This can be visualizaed by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example). The values of $R$, $G$ and $B$ are determined by looking at the image of said number at each stage of the iteration, $N_{p}^{n}(z_{0})$, and computing the current distance to each of the roots of $p(z)$; so $R=1/d_{r}$ where $d_{r}$ is the positive distance to the root which is colored red, and so on. In this way, the roots themselves are colored red, green, and blue, and every other point gets a mix of the three colors. As the number of iterations increases, each point gets a sharper color, as the sequence of images $\{N_{p}^{n}(z_{0})\}_{n=1}^{\infty}$ converges to one of the three roots. Each stage, the complex plane looks as if out of focus because the coloring function is continuous. As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton fractal (of which, interestingly, Newton was unaware).
@@
We can frame Newton’s method as a CCS (satisfying the Extendibility Axiom) as follows. Given a non-constant polynomial $p(z)$, let $L\subseteq\mathbb{C}$ be the complement of the Julia set (this is called the Fatou set) associated to $N_{p}$. This is the set where Newton’s method is tame. By [BLA84, Corollary 4.6], $L$ is open and dense in $\mathbb{C}$. Let $\mathcal{P}=\{P_{1},P_{2},P_{3}\}$ where the map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$, i.e.,
$P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
$P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
$P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
Since $L$ is dense in the extended complex plane, then $\mathcal{L}=S^{2}$. Computations are the iterations of Newton’s method, i.e., $\Delta=\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$. Then, $(L,\mathcal{P},\Gamma)$ is a CCS. Since all computations $N_{p}^{n}$ are rational maps, they can be continuously extended to the extended complex plane $\overline{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, i.e., to $\mathcal{L}$. In particular, $(L,\mathcal{P},\Gamma)$ satisfy the Extendibility Axiom. The set of deep computations $\overline{\Delta}$ might behave different for various polynomials. Let us look at various examples:
@@
The Newton’s method map $N_{p}:L\rightarrow L$ is given by
@@
Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$. For example,
@@
However, $\tilde{f}$$\mathcal{L}\rightarrow\mathcal{L}$ is not continuous. The set $\overline{\tilde{\Delta}}$ of deep computations is $\tilde{\Delta}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
@@
Newton’s method for $p(z)=z^{3}-2z+2$. Let $r_{1}$, $r_{2}$ and $r_{3}$ be the three roots of $p(z)$. Let $B_{1}$, $B_{2}$ and $B_{3}$ be their respective basins. Let $B$ be the basin of the attractive cycle $0,1,0,1,\dots$. Let $L=B\cup\bigcup_{i=1}^{3}B_{i}$ and $J$ be its complement, i.e., the Julia set. Notice that $N_{p}^{n}$ does not converge pointwise. However, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to functions $f_{1}$ and $f_{2}$ respectively. $f_{1}$ and $f_{2}$ are two distinct deep computations. Note that for $z\in J$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number. However, since $\mathcal{L}=S^{2}$ is compact there is a subsequence of $\tilde{N_{p}^{n}}(z)$ that converges to $\infty$. We can extend $f_{i}:L\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
$\tilde{f_{i}}(z)=\begin{cases}f_{i}(z),&\text{if }z\in L;\\ \infty,&\text{if }z\in J.\end{cases}.$
Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Delta}}$.
@@ 2.4. Finite precision threshold classifiers as a CCS
Let $L=2^{\mathbb{N}}$, i.e., the set consisting of all infinite binary sequences with the topology of pointwise convergence. Let $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ is the collection of projections, i.e., $P_{n}(x)=x(n)$ for each $n$ and each $x\in L$. In other words, $P_{n}$ is the predicate that reads the $n$th digit of a sequence. Notice that $L\subseteq\mathbb{R}^{\mathcal{P}}$ is closed. Therefore, $\mathcal{L}=L$. We denote by $0^{\infty}$ the infinite binary sequence consisting of 0s, and by $1^{\infty}$ the infinite binary sequence consisting of $1$s. The set of finite binary strings is denoted by $2^{<\mathbb{N}}$. Given a finite binary string $w$, we consider the transition $\phi_{w}:L\to L$ given by the rule
$\phi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0^{\infty},&\text{otherwise};\end{cases}$
@@
We first prove that $f^{-1}(1)=\{x\in 2^{\mathbb{N}}:x|_{|w|}\leq_{\mathrm{lex}}w\}$ is an open set. Fix $x_{0}\in f^{-1}(1)$. Let $t:=x_{0}|_{|w|}$ and consider the basic open set $t]=\{x\in 2^{\mathbb{N}}:x|_{|t|}=t\}$. Then, it is not difficult to check that $x_{0}\in[t]\subseteq f^{-1}(1)$. The same reasoning shows that $f^{-1}(0)$ is open. ∎
Let $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$ where $\mathbf{0^{\infty}},\mathbf{1^{\infty}}:L\rightarrow L$ are the constant maps identical to $0^{\infty}$ and $1^{\infty}$, respectively. Let $\Gamma$ be the semigroup generated by $\Delta$. The preceding proposition shows that $\Delta$ (and hence $\Gamma$) consists of continuous functions. In particular, the CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom. In contrast with Newton’s method, the algebraic structure of $\Delta$ is quite simple: composing two classifiers results in something similar to a Boolean logic gate. The topological structure is far more interesting. Intuitively, the crucial difference between Newton’s method and threshold classifiers is that the complexity of the former comes from depth: the semigroup is generated by a single map but its iterates are highly complex. The complexity of threshold classification comes from width: the semigroup has infinitely many generators, but their compositions are simple.
Intuitively, the closure of $\Delta$ consists of the set of all possible threshold classifiers on the real line, but there are two sorts: the ones that classify strict inequalities and those that classify $\leq$. The members of $\Delta$ are finite-precision approximations of classifiers that check all bits of information. But here it gets interesting: what is the difference, in terms of floating-point arithmetic, between $x<0.5$ and $x\leq 0.5$?
@@
The preceding proposition shows that the topological structure of deep computations can be very complicated. Indeed, $\overline{P_{n}\circ\Delta}$ contains the Split Cantor space for all $n\in\mathbb{N}$. (see Examples 3.3(3)).

### v3

@@ A Compositional Computation Structure (CCS) is a triple $(L,\mathcal{P},\Gamma)$, where
For a CCS $(L,\mathcal{P},\Gamma)$, we regard the elements of $\Gamma$ as “standard” finitary computations, and the elements of $\overline{\Gamma}$, i.e., deep computations, as possibly infinitary limits of standard computations. The main goal of this paper is to study the computability, definability and computational complexity of deep computations. Since deep computations are defined through a combination of topological concepts (namely, topological closure) and structural and model-theoretic concepts (namely, models and types), we will import technology from both topology and model theory.
@@ 2.2. Computability and definability of deep computations and the Extendibility Axiom
    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{R}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
@@
    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]). The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
@@
To summarize, a function $f:L\to\mathcal{L}$ is computable if and only if it is definable if and only if it is polynomially approximable if and only it can be extended to a continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$. This is the reason for the following definition.
@@
For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$. For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.
@@ 2.3. Newton’s method as a CCS
Let $p(z)$ be a non-constant polynomial with complex coefficients. We say that $(L_{p},\mathcal{P},\Gamma_{p})$ is Newton’s method for $p(z)$ if:
  • $L_{p}$ is the set of all $z\in\mathbb{C}$ such that there exists an open neighborhood $U$ of $z$ such that every sequence in $\{N_{p}^{n}:n\in\mathbb{N}\}$ has a subsequence that converges uniformly on compact subsets of $U$, where
@@
    and $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ is the nth-iteration of $N_{p}$.
  • $\mathcal{P}:=\{P_{1},P_{2},P_{3}\}$ where
    $P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
    $P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
    $P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
  • $\Gamma_{p}:=\{N_{p}^{n}:n\in\mathbb{N}\}$.
Remarks 2.6.
  1. (1)
    The set $L_{p}$ is known as the Fatou set of $N_{p}$ (see section 2 in [BLA84]). Its complement $\mathbb{C}\backslash L_{p}$, is known as the Julia set of $N_{p}$.
  2. (2)
    The map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$.
  3. (3)
    The set $L_{p}$ is open and dense in $\mathbb{C}$ ([BLA84], Corollary 4.6). Hence, its closure $\mathcal{L}$ in $\mathbb{R}^{\mathcal{P}}$ is the Riemann sphere $S^{2}$ (i.e., the extended complex plane).
  4. (4)
    The set $L_{p}$ is completely invariant under iterations of $N_{p}$, i.e., $N_{p}(L_{p})=L_{p}$ and $N_{p}^{-1}(L_{p})=L_{p}$. This implies that all iterations $N_{p}^{n}:L_{p}\rightarrow L_{p}$ are transition maps.
  5. (5)
    $\Gamma_{p}$ is the semigroup generated by $\{N_{p}\}$. Thus, $(L_{p},\mathcal{P},\Gamma_{p})$ is a CCS.
Newton’s method is an iterative method that is used to approximate a root of $p(z)$. The map $N_{p}(z)$ defined above is known as Newton’s map. The method consists of taking an initial guess $z_{0}\in\mathbb{C}$ and iterating the rational map $N_{p}$ to obtain a sequence given by
@@
For each root $r\in\mathbb{C}$ of $p(z)$, there exists an $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
Given a root $r$ of $p(z)$, the set $B_{r}=\{z\in\mathbb{C}:\lim_{n\rightarrow\infty}N_{p}^{n}(z)=r\}$ is an open set called the basin of $r$. However, Newton’s method can fail to converge to any root for some choices of $z_{0}$. For example, consider the polynomial $p(z)=z^{3}-2z+2$. The Newton map is given by
@@
Notice that taking $z_{0}=0$ as an initial guess will yield the sequence $0,1,0,1,0,1,\dots$ that oscillates between 0 and $1$ but none of them are roots of $p(z)$. Another more chaotic way Newton’s method can fail to converge is when the sequence of iterations has no convergent subsequence. The set of such points, i.e. the Julia set associated to $N_{p}$, is typically a fractal. This can be visualized by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example). The values of $R$, $G$ and $B$ are determined by looking at the image of said number at each stage of the iteration, $N_{p}^{n}(z_{0})$, and computing the current distance to each of the roots of $p(z)$; so $R=1/d_{r}$ where $d_{r}$ is the positive distance to the root which is colored red, and so on. In this way, the roots themselves are colored red, green, and blue, and every other point gets a mix of the three colors. As the number of iterations increases, each point gets a sharper color, as the sequence of images $\{N_{p}^{n}(z_{0})\}_{n=1}^{\infty}$ converges to one of the three roots. At each stage, the complex plane looks as if out of focus because the coloring function is continuous. As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton’s fractal (of which, interestingly, Newton was unaware).
@@
Proposition 2.7.
Let $p(z)$ be a non-constant polynomial. $(L_{p},\mathcal{P},\Gamma_{p})$ satisfies the Extendibility Axiom.
Proof.
$N_{p}:L_{p}\rightarrow L_{p}$ is a rational map. Rational maps can be continuously extended to the extended complex plane, i.e., to $\mathcal{L}$. Composition of rational maps is a rational map, so by the same reasoning, computations $N_{p}^{n}:L_{p}\rightarrow L_{p}$ can be continuously extended to $\mathcal{L}$. ∎
The set of deep computations $\overline{\Gamma}$ might behave different for various polynomials. Let us look at various examples:
@@
Let $\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$ where
@@
As before, $(L,\mathcal{P},\Gamma)$ is a CCS. Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$. For example,
@@
However, $\tilde{f}$$\mathcal{L}\rightarrow\mathcal{L}$ is not continuous. The set $\overline{\tilde{\Gamma}}$ of deep computations is $\tilde{\Gamma}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
@@
Newton’s method for $p(z)=z^{3}-2z+2$. Let $r_{1}$, $r_{2}$ and $r_{3}$ be the three roots of $p(z)$. Let $B_{1}$, $B_{2}$ and $B_{3}$ be their respective basins. Let $B$ be the basin of the attractive cycle $0,1,0,1,\dots$. Then, $L_{p}=B\cup\bigcup_{i=1}^{3}B_{i}$. Notice that $N_{p}^{n}$ does not converge pointwise. However, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to functions $f_{1}$ and $f_{2}$ respectively. $f_{1}$ and $f_{2}$ are two distinct deep computations. Note that for $z\notin L_{p}$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number. However, since $\mathcal{L}=S^{2}$ is compact there is a subsequence of $\tilde{N_{p}^{n}}(z)$ that converges to $\infty$. We can extend $f_{i}:L_{p}\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
[…]
Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Gamma}}$.
@@ 2.4. Finite precision threshold classifiers as a CCS
Let $L=2^{\mathbb{N}}$, i.e., the set consisting of all infinite binary sequences with the topology of pointwise convergence. Let $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ equal the collection of projections, i.e., $P_{n}(x)=x(n)$ for each each $x\in L$ and $n\in\mathbb{N}$. Notice that $L\subseteq\mathbb{R}^{\mathcal{P}}$ is closed. Therefore, $\mathcal{L}=L$. We denote by $0^{\infty}$ the infinite binary sequence consisting of 0s, and by $1^{\infty}$ the infinite binary sequence consisting of $1$s. The set of finite binary strings is denoted by $2^{<\mathbb{N}}$. This set is naturally ordered by the lexicographic order $\leq_{\mathrm{lex}}$. Given a finite binary string $w$, we consider the transition $\phi_{w}:L\to L$ given by the rule
$\phi_{w}(x)=\begin{cases}1^{\infty},&\text{if }x|_{|w|}\leq_{\mathrm{lex}}w;\\ 0^{\infty},&\text{otherwise},\end{cases}$
@@
We first observe that $f^{-1}(1)=\{x\in 2^{\mathbb{N}}:x|_{|w|}\leq_{\mathrm{lex}}w\}$ is an open set. Fix $x_{0}\in f^{-1}(1)$. Let $t:=x_{0}|_{|w|}$ and consider the basic open set $t]=\{x\in 2^{\mathbb{N}}:x|_{|t|}=t\}$. Then it is not difficult to check that $x_{0}\in[t]\subseteq f^{-1}(1)$. The same reasoning shows that $f^{-1}(0)$ is open. ∎
Let $\Delta=\{\phi_{w}:w\in 2^{<\mathbb{N}}\}\cup\{\mathbf{0^{\infty}},\mathbf{1^{\infty}}\}$, where $\mathbf{0^{\infty}},\mathbf{1^{\infty}}:L\rightarrow L$ are the constant maps identical to $0^{\infty}$ and $1^{\infty}$, respectively. Let $\Gamma$ be the semigroup generated by $\Delta$. The preceding proposition shows that $\Delta$ (and hence $\Gamma$) consists of continuous functions. In particular, the CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom. In contrast with Newton’s method, the algebraic structure of $\Delta$ is quite simple: composing two classifiers results in something similar to a Boolean logic gate. The topological structure is far more interesting. Intuitively, the crucial difference between Newton’s method and threshold classifiers is that the complexity of the former comes from depth: the semigroup is generated by a single map but its iterates are highly complex. The complexity of threshold classification comes from width: the semigroup has infinitely many generators, but their compositions are simple.
Intuitively, the closure of $\Delta$ consists of the set of all possible threshold classifiers on the real line, but there are two sorts: the ones that classify strict inequalities and those that classify $\leq$. The members of $\Delta$ are finite-precision approximations of classifiers that check all bits of information. But here it gets interesting: what is the difference, in terms of arbitrary-precision arithmetic, between $x<0.5$ and $x\leq 0.5$?
@@
The preceding proposition shows that the topological structure of deep computations can be complicated. Indeed, $\overline{P_{n}\circ\Delta}$ contains the Split Cantor space for all $n\in\mathbb{N}$. (see Examples 3.3(3)).